Ok, I have tried to solve this equation :
Find f(a+2) for the function f(x)=3x^2-x+2
and I keep getting 3a^2-a+4, can someone please show me how to solve this equation?
Thank you.

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

Ok, I have tried to solve this equation :
Find f(a+2) for the function f(x)=3x^2-x+2
and I keep getting 3a^2-a+4, can someone please show me how to solve this equation?
--------------------
Sub a+2 for x
f(a+2) = 3(a+2)^2 - (a+2) + 2
= 3a^2 + 12a + 12 -a-2 + 2
= 3a^2 + 11a + 12

Answer by solver91311(24713)


Whatever is in the parentheses after the function designator is called the argument of the function. Always substitute this entire argument for the independent variable in the function definition.

So if



then







and finally,



but





John